When 25 g of ice at -20°c have absorbed 2500 cal of heat, what os the temperature of the resulting water?


1
Expert's answer
2021-05-10T02:42:19-0400

Heat required by ice to melt to water

Q=mL=25*80=2000calories

Heat given is 2500 cal

Heat given is more than required thus all ice melt and then temperature rises.

Heat responsible to rise the temperature of water=2500-200=500cal

Q=mcdt

500=25*1*(Tf-0)

TF=500/25=20 degree Centigrade